OpenStudy (anonymous):

At a skills competition, a target is being lifted into the air by a cable at a constant speed. An archer standing on the ground launches an arrow toward the target. The system of equations below models the height, in feet, of the target and the arrow t seconds after it was fired. Which statement most likely describes the situation modeled by this system? mc016-1.jpg The arrow is fired with an initial upward velocity of 2 ft/s. The arrow is fired with an initial upward velocity of 4 ft/s. The arrow is fired with an initial upward velocity of 8 ft/s. The arrow is fired with an initial upw

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

I can't see the equations.

OpenStudy (dumbcow):

The height of the arrow is modeled by \[h(t) = -16t^2 + vt + h_0\] The height of the target is \[H(t) = st + H_0\] Set 2 equations equal to find velocity v needed to hit the target (assuming you are given time t)

OpenStudy (dumbcow):

you to repost the link so we can see the diagram

OpenStudy (anonymous):

h= 8 + 2t h = 4 + 32t-16tsquared